OpenStudy (anonymous):

Let ABC a triangle. Take n points on the side AB, join them with C by straight lines, take n points on BC, join them with A by straight lines, take n points on AC, join them with B by straight lines. If no three straight lines meet at any point except at A,B,C then find the number of partitions in the triangle. Please show the solution with proof

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

(1) When you draw a line from a vertex to the opposite side, each time you meet another line you create one more region. Initially there is 1 region (the original triangle) When you draw a line from C to AB, you meet one line (the side AC) so each line creates one more region. n lines create n+1 regions altogether (so far, the result is obvious) (2) When you draw a line from B to AC, you meet n+1 lines (the n lines you drew in step 1 and the side AC. So each line create another n+1 regions Adding a lines create n(n+1) regions, plus the original n+1, = (n+1)^2 (3) When you draw a line from A to BC you meet 2n+1 lines (the n lines from B, the n lines form C, and the side BC) Total number of regions = n(2n+1) + (n+1)^2 = 3n*2 + 3n + 1 http://in.answers.yahoo.com/question/index?qid=20100210065419AAbX9uq

OpenStudy (anonymous):

@Arnab09 do u get it

OpenStudy (anonymous):

I am not sure about the first statement. If it is true, then its alright :)

OpenStudy (anonymous):

sorry, I posted this problem in physics section by mistake :(